q1= 20nC, q2=10nC located on the x axis. x=0 and x=1 meter respectively. Where on the x axis is electric field zero?
1
Expert's answer
2020-01-31T10:40:31-0500
"E=\frac{q_1}{x-0}-\frac{q_2}{1-x}=0"

"\frac{20}{x-0}=\frac{10}{1-x}"

"x=\frac{2}{3}m=0.67\ m"